OpenStudy (anonymous):

make x the subject: y = x+2/x+3

OpenStudy (anonymous):

You mean isolate the x, right?

OpenStudy (anonymous):

yeah, i have no idea how to start doing it

OpenStudy (anonymous):

Oh, I see, never seen someone writes "make x the subject." Multiply both sides by x+3\[y(x+3) = x+2\]\[xy + 3y = x+ 2\]Now take all terms with x to one side while take all terms that don't have x to the other side.\[xy-x = 2-3y\]\[x(y-1)= 2-3y\]Divide both sides by y-1.\[\boxed{x = \dfrac{2-3y}{y-1}}\] Is this clear?

OpenStudy (anonymous):

i dont understand how you got 2 -3y?

OpenStudy (anonymous):

oh wait dont worry i get it haha

OpenStudy (anonymous):

I subtracted both sides by 3y.

OpenStudy (anonymous):

oh ok haha

OpenStudy (anonymous):

yeah took a while to see that lol

OpenStudy (anonymous):

thanks for the help mate

OpenStudy (anonymous):

No problem.